HI. PLEASE HELP ME WITH THIS PROBLEM : IF f(4)=0 and f(6)=6, which of the following could represent f(x)? THANK YOU!
A. 2/3X-4
B. X+2
C. X-4
D. 3/2X+6
E. 3X-12

Answer by scott8148(6628)

you have to substitute 4 and 6 into the functions, and see if you get 0 and 6 as results

A. (2/3)(4)-4<0 ... NO

B. (4)+2>0 ... NO

C. (4)-4=0, (6)-4<6 ... NO

D. (3/2)(4)+6>0 ... NO

E. 3(4)-12=0, 3(6)-12=6 ... WINNER!!